OpenStudy (romantic_ch3micals):

Medal. Please explain and show work because idk what to do. Steve is driving to a ski resort. At first, he drove at a rate of 65 mph. Then a snow storm started and he had to slow down to 40 mph. In total. it took him 6 hours to drive 310 miles. Which system of equations represents this situation? A. 6x + y= 1 x + y= 310 B. x + y= 105 6x + 40y=310 C. 310x + 65y= 40 x + y= 65 D. x + y=6 65x + 40y=310

OpenStudy (anonymous):

idk the answer but chemical romance is life yo

jagr2713 (jagr2713):

if he drove at rate 60 mph , then x descripe the time he travel with this speed; y descripte the time he travel after the storm

jagr2713 (jagr2713):

60x the distance d=time * speed

jagr2713 (jagr2713):

40y distance

jagr2713 (jagr2713):

6x+40y = there all distance 310 miles according to given data; x+y is the speed of them = 65+40 =105

OpenStudy (romantic_ch3micals):

thanks

jagr2713 (jagr2713):

YW

OpenStudy (romantic_ch3micals):

wait...so it's B

OpenStudy (anonymous):

Yeah its B. @Romantic_Ch3micals

OpenStudy (romantic_ch3micals):

ok